### v1

# DynFocus: Dynamic Cooperative Network Empowers LLMs with Video Understanding

Abstract
The challenge in LLM-based video understanding lies in preserving visual and semantic information in long videos while maintaining a memory-affordable token count. However, redundancy and correspondence in videos have hindered the performance potential of existing methods. Through statistical learning on current datasets, we observe that redundancy occurs in both repeated and answer-irrelevant frames, and the corresponding frames vary with different questions. This suggests the possibility of adopting dynamic encoding to balance detailed video information preservation with token budget reduction. To this end, we propose a dynamic cooperative network, DynFocus, for memory-efficient video encoding in this paper. Specifically, i) a Dynamic Event Prototype Estimation (DPE) module to dynamically select meaningful frames for question answering; (ii) a Compact Cooperative Encoding (CCE) module that encodes meaningful frames with detailed visual appearance and the remaining frames with sketchy perception separately. We evaluate our method on five publicly available benchmarks, and experimental results consistently demonstrate that our method achieves competitive performance.
((a))
((b))
Figure 1: Concept of redundancy and correspondence in our pipeline. (a) The proportion of redundancy for video datasets [^2] The implementation is provided in supplementary materials.. Redundancy includes both repeated and answer-irrelevant frames. Repeatance gauges the redundancy between consecutive frames, while answer-irrelevance refers to frames with a marginal contribution to question answering. (b) An example of correspondence. Given a video, we highlight the corresponding question/answer pairs and frames using red and blue boxes, respectively.
[…]
1 Introduction
Large Language Models (LLMs) have shown their ability on general AI  [22]. Vision Language Models (VLMs) extend the capabilities of LLMs to process visual data, demonstrating proficiency in tasks such as image captioning and visual question answering. However, challenges arise in video understanding, especially with long-term videos, where representing consecutive video frames requires an excessive number of tokens, leading to high memory usage.
Recent attempts use average pooling, attention, or dynamic masking to reduce video tokens spatially [28, 37, 56, 21, 35]. However, redundant frames lead to the neglect of key visual details. Several works [15, 55] capture visual appearance with memory banks to preserve key details. However, the key details vary in correspondence to questions, which can easily result in the loss of keyframes from long videos and increase the overhead of the memory banks.
In Fig. 1, we illustrate examples of redundancy and correspondence in videos. We observe that i) there is significant redundancy among frames, with only a few meaningful frames directly contributing to question answering. This suggests the potential for adopting a dynamic frame encoding strategy to reduce tokens based on their contribution. ii) Answering different questions generally requires focusing on different parts of the frame. Therefore, dynamically identifying meaningful frames offers better flexibility for sophisticated video content understanding.
In this paper, we propose a dynamic cooperative framework, DynFocus, for memory-efficient video encoding. Specifically, it consists of two key modules: Dynamic Event Prototype Estimation (DPE) and Compact Cooperative Encoding (CCE). DPE serves as the dynamic selector to accurately discern the meaningful frames, which takes both redundancy and correspondence to question answering into consideration as the selection standard. Afterwards, CCE complies with the dynamic encoding principle. The meaningful frames are encoded with fine-grained context features for detailed visual appearance, whereas those redundant frames are substantially encapsulated into a few tokens for sketchy perception, which enables LLM to capture broader temporal clues within a fixed receptive field. These two modules reconcile the nuanced visual-semantic understanding with affordable token quota.
Figure 2: Competitive even SOTA performance on various video understanding tasks.
Moreover, our CCE module draws inspiration from the cooperation of retinal ganglion cells in the primate visual system. Biological studies [18, 45] found that in these cells, Rod cells perceive the overall scene in a wide field of view, while Cone cells understand complex scenes with fine details. These cells are located at the periphery of the retina and arranged in a parallel manner, receiving the signals but activated under different conditions. Our framework is analogous in two aspects: (1) Which cell is activated depends on whether the current input frame is meaningful or not. (2) The meaningful frames are encoded with fine-grained tokens as key detailed clues, akin to Cones, whereas the marginal frames are condensed into low-resolution tokens, ensuring better temporal consistency, similar to Rods. We hope these relations will further support the design philosophy of our method and reveal its rationality.
Our contributions are summarized as follows:
  • We propose a dynamic cooperative network, DynFocus, towards memory-efficient video encoding within LLM, inspired by the biological concept of Cone and Rod cells.
  • We introduce two modules, DPE and CCE, that dynamically balance subtle visual appearance with sketchy temporal perception using affordable tokens.
  • Experimentally, we evaluate our DynFocus on two publicly mainstream short video benchmarks, three long video benchmarks, and one diagnosis benchmark on video hallucination as shown in Fig. 2.
2 Related Work
Video-based Large Language Models. In recent years, Vision Language Models (VLMs) has emerged to extend the capabilities of LLMs [42, 8, 52, 7, 40] to handle diverse and complicated inputs with satisfactory generalization. Generally, VLMs incorporate additional connector to bridge the semantic gap between input video content and LLMs, further performing modality alignment and instruction tuning on video-based dataset. However, video understanding presents the significant challenges due to their extensive memory overhead. Several studies have dedicated to addressing these challenges with greater efficiency. Video-ChatGPT [37] adopts both spatial and temporal pooling to condense video tokens. VideoChat [25] employs a learnable Q-former [9] to aggregate the similar tokens for memory reduction. Chat-UniVi [21] develops a unfied framework for processing both image and video, which reduces spatial and temporal tokens through multi-stage token merging. Although these methods alleviate the memory usage to some extent, they often discard the abundant temporal clues by sampling parts of frames as the input. To compensate the loss of temporal clues, LLaMA-VID [28] innovates with a dual-token approach that represents each frame with context and content tokens, which allows for larger video throughput. MovieChat [50] incorporates the short-term memory and long-term memory into unified framework, strategically combining similar frames to reduce memory footprint while capture the temporal clues. Similarly, MA-LLM [16] stores past video information in a memory bank, which allows to reference historical video content for long-term analysis without exceeding memory limits. However, these methods exhibit proficiency in capturing temporal clues at the expense of discarding the visual details. In a nutshell, they struggle to jointly capture the spatial details and temporal dynamics effectively.
Dynamic Networks. Dynamic networks, adjusting the encoding strategy according to specific input, have recently garnered the burgeoning interest across various domains. Early methods mainly focus on traditional image classification by channel pruning or layer skipping. For example, BlockDrop [57] designed an auxiliary policy network to determine whether skip or execute convolutional blocks via reinforcement optimization. Based on dynamic mechanism, a series of research efforts are devoted to better adapting to the various dynamic scenes. Specifically, Dynamic [27] proposes a routing network with soft conditional gate to adaptively search data-dependent scale transformation paths for semantic segmentation. In the field of image question answering, SUPER [14] develops a semantic-aware modular routing framework to recursively handle different complexity of visual scene. In this work, we marks the first attempt to reveal the substantial potential of dynamic encoding strategies when understanding the complicated long-term video.
Figure 3: Schematic Illustration of DynFocus. Our method takes the user instruction and video frames as input, and yields the compact video tokens from CCE module for LLM. Specifically, DPE module serves as the selector to identify the prototypes that contribute greatly to answer, providing CCE module with event prototype $\{\mathbf{h}_{k}\}_{k=1}^{K}$ and the binary mask $\{b_{t}\}_{t=1}^{T}$, which is marked with two red arrows. Benefited from this, CCE module dynamically encode the critical prototypes with more tokens, and encapsulate the marginal prototypes with few tokens. T-DPC and S-DPC represent the DPC-KNN clustering temporally and spatially, respectively.
3 Dynamic Cooperative Network
As shown in Fig 3, the overall framework is comprised of three parts. Part 1: visual and text encoder are adopted to produce the corresponding features; Part 2: the proposed dynamic cooperative network serves as the connector to compress the video content for LLM, which consists of two modules, DPE module and CCE module; Part 3: the foundational LLM receives the token sequence outputted from Part 2 to generate the language response.
3.1 Visual and Text Encoder
Given a $T$-frame video, we extract the frame-wise features $\mathbf{V}=\{\mathbf{v}_{t}\}_{t=1}^{T}$, using a pre-trained visual encoder, where $\mathbf{v}_{t}\in\mathbb{R}^{N\times d}$ denotes the feature of the $t$-th frame. Here, $N$ is the number of image patches and $d$ is the feature dimension. For the text encoder, user instruction is fed to the pre-trained text encoder to generate the text features $\mathbf{Q}=\{\mathbf{q}_{r}\}_{r=1}^{R}$, where $\mathbf{q}_{r}\in\mathbb{R}^{d^{\prime}}$ denotes the feature embedding of each word in user instruction. $d^{\prime}$ is the feature dimension and $R$ is the total number of words.
3.2 Dynamic Event Prototype Estimation
Given frame-wise features $\mathbf{V}$, we aim to estimate event prototypes, i.e., discriminative features that are most relevant to ground-truth answer.
For each frame feature $\mathbf{v}_{t}$, we first perform local average pooling spatially. The pooling process reduces the number of features from $N$ to $P$ for efficiency, resulting in $\mathbf{f}_{t}=\mathrm{Pool}(\mathbf{v}_{t}),\mathbf{f}_{t}\in\mathbb{R}^{P\times d}$. Redundancy exists in $T$-frame features $\mathbf{F}=\{\mathbf{f}_{t}\}_{t=1}^{T}$. To remove redundancy and identify representative features from $\mathbf{F}$, we perform clustering on $\mathbf{F}$, and use cluster centers to estimate event prototypes.
Following [21], we borrow components from the traditional DPC-KNN [11] algorithm for obtaining cluster centers. For self-contain purposes, we briefly summarize two important variables (local density $\rho_{t}$ and distance indicator $\delta_{t}$ ) in DPC-KNN below.
For clustering, the local density $\rho_{t}$ measures the mean distance to $K$ nearest neighbors of the $t$-th frame, and is given by,
$\rho_{t}=\mathrm{exp}\left(-\frac{1}{K}\sum_{{t^{\prime}}\in\mathcal{N}({t})}\frac{1}{P}\|\mathbf{f}_{t}-\mathbf{f}_{t^{\prime}}\|_{F}^{2}\right),$ (1)
where ${{t^{\prime}}\in\mathcal{N}({t})}$ denotes that $\mathbf{f}_{t^{\prime}}$ is in the neighborhood of $\mathbf{f}_{t}$. $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm, and it is also used to perform $K$ nearest neighbor search for each $\mathbf{f}_{t}$.
The distance indicator $\delta_{t}$ measures the possibility of $t$-th frame to be cluster center by calculating the minimum distance between $\mathbf{f}_{t}$ and any other frames with higher density, and is given by,
$\delta_{t}=\left\{\begin{array}[]{l}\underset{t^{\prime}}{\min}\frac{1}{P}\left\|\mathbf{f}_{t}-\mathbf{f}_{t^{\prime}}\right\|^{2}_{F},\text{ if }\exists t^{\prime}\text{ s.t. }\rho_{t^{\prime}}\textgreater\rho_{t}\\ \underset{t^{\prime}}{\max}\frac{1}{P}\left\|\mathbf{f}_{t}-\mathbf{f}_{t^{\prime}}\right\|^{2}_{F},\text{ otherwise }.\end{array}\right.$ (4)
We use the product of $\rho_{t}$ and $\delta_{t}$ to measure the importance of each frame. Frames with high scores are more likely to be informative. We first sort scores $\{\rho_{t}\times\delta_{t}\}_{t=1}^{T}$ in the decreasing order and then take the Top-$L$ frame features from $\mathbf{F}$ with high scores as representative features. Finally, we normalize representative features using their importance scores and estimate event prototypes as follows,
$\mathbf{m}_{l}=\frac{\sum_{t^{\prime}\in\mathcal{N}(l)}\mathrm{exp}(\rho_{t^{\prime}}\cdot\delta_{t^{\prime}})\mathbf{f}_{t^{\prime}}}{\sum_{t^{\prime}\in\mathcal{N}(l)}\mathrm{exp}(\rho_{t^{\prime}}\cdot\delta_{t^{\prime}})},l\in[1,L],$ (5)
where $\mathrm{exp}(\rho_{t^{\prime}}\cdot\delta_{t^{\prime}})$ denotes the importance weight of $\mathbf{f}_{t^{\prime}}$.
Note that the obtained event prototypes $\mathbf{M}=\{\mathbf{m}_{l}\}_{l=1}^{L}$ from Eq. (5) are estimated only based on frame features, and are not aligned to the ground-truth answer. In other words, $\mathbf{m}_{l}\in\mathbb{R}^{P\times d}$ may contain visual redundancy, which would disturb useful clues for sophisticated video understanding. In light of this, we aim to further select answer-relevant event prototypes from $\mathbf{M}$. Specifically, we use a multi-layer perceptron (MLP) network $\mathcal{U}(\cdot)$ to regress frame-wise scores. Note that $\mathcal{U}(\cdot)$ is learned end-to-end with the supervision from LLM, thus is aligned to ground-truth answer implicitly. The regression is given by,
$s_{l}=\mathcal{U}\left(\mathrm{Max}(\mathbf{m}_{l})||\mathrm{Avg}(\mathbf{m}_{l})\right),$ (6)
where $\mathrm{Max}(\mathbf{m}_{l})\in\mathbb{R}^{d}$ denotes the row feature with maximum feature norm ($L_{2}$) across $P$ rows. $\mathrm{Avg}(\mathbf{m}_{l})\in\mathbb{R}^{d}$ denotes the averaged row feature across $P$ rows. |$\cdot$|$\cdot$ | | ⋅ denotes concatenation. After collecting all scores into a score vector $\mathbf{s}=\{\mathbf{s}_{l}\}_{l=1}^{L}$, we perform min-max normalization to normalize score values to be within $[0,1]$. Finally, we sort scores in the decreasing order and take the Top-$K$ event prototypes from $\mathbf{M}$ with high scores as filtered event prototypes, denoted by $\mathbf{H}=\{\mathbf{h}_{k}\}_{k=1}^{K}$. We also retrieve indices of filtered event prototypes in the $T$-frame video, obtaining the binary index mask $\mathbf{b}=\{b_{t}\}_{t=1}^{T}$. Notably, $b_{t}=1$ indicates important frame, while $b_{t}=0$ signifies redundant frame.
Training $\mathcal{U}(\cdot)$. Note that the Top-$K$ operation is not differentiable and thus stops the gradient propagation from LLM to update our score net $\mathcal{U}(\cdot)$. This limitation restricts our $\mathcal{U}(\cdot)$ to dynamically estimate the redundancy and flexibly capture correspondence without auxiliary loss supervision. To address this issue, we transform the Top-$K$ operation into solving a linear programming problem to make our network end-to-end trainable, and it can be directly updated by the autoregressive loss between language response and ground-truth answer within LLM. Please refer to the supplementary material for details.
3.3 Compact Cooperative Encoding
Given frame-wise features $\mathbf{V}=\{\mathbf{v}_{t}\}_{t=1}^{T}$, event prototypes $\mathbf{H}=\{\mathbf{h}_{k}\}_{k=1}^{K}$, and their corresponding index mask $\mathbf{b}=\{b_{t}\}_{t=1}^{T}$, we perform cooperative encoding for memory-efficient video understanding. Frames that contribute greatly to answer (i.e., $b_{t}=1$) will be encoded with more tokens than marginal frames (i.e., $b_{t}=0$), to capture intricate spatial details.
For each frame feature $\mathbf{v}_{t}$, we use the same clustering pipeline (Eq. (1),(4),(5)) in Sec. 3.2 to estimate spatial object prototypes $\mathbf{Z}_{t}=\{\mathbf{z}_{t,i}\}_{i=1}^{I}$. The only difference is that spatial clustering is performed on $N$ patch features within single frame feature $\mathbf{v}_{t}$, aggregating $N$ patch features into $I$ prototypes. In contrast, temporal clustering in Sec. 3.2 is performed on $T$ frame features in the whole video.
We find that semantic abstraction of $\mathbf{v}_{t}$ can be achieved when spatial clustering is performed multiple times. For example, concepts such as “person” and “dog” are progressively formed from the low-level attribute or color information. We thus use multiple clustering layers to capture more abundant visual details. The output spatial object prototypes of layer $j$ is fed to layer $j+1$ for abstraction. Collecting all layer outputs results in our multi-grained spatial object prototypes $\mathbf{G}_{t}=\{\mathbf{Z}^{(j)}_{t}\}_{j=1}^{J}$. Here, $J$ is the total number of clustering layers.
Cones Encoding.
We mimic cones to focus on fine visual appearance. Specifically, frames with $b_{t}=1$ are encoded with the combination of event and its corresponding multi-grained spatial prototypes,
$\mathbf{U}_{t,b_{t}=1}=\mathcal{F}_{\text{fine}}\left(\mathbf{h}_{t}||\mathbf{G}_{t}\right),$ (7)
where $\mathcal{F}_{\text{fine}}$ is a simple MLP network. Note that no feature pooling operation is performed to capture delicate details, and the number of tokens in $\mathbf{U}_{t,b_{t}=1}$ equals the number of summation of event and multi-grained spatial prototypes.
Rods Encoding.
We mimic rods to focus on coarse temporal dynamics towards broader video understanding. Specifically, frames with $b_{t}=0$ are encoded with the modulation of text embedding $\mathbf{Q}$, to obtain text-grounded visual clues,
$\mathbf{E}=\mathrm{Softmax}\Bigg{(}\frac{f_{q}(\mathbf{G}_{t})(f_{k}(\mathbf{Q}))^{\top}}{\sqrt{d}}\Bigg{)}\mathbf{G}_{t},$ (8)
where $f_{q}(\cdot)$ and $f_{k}(\cdot)$ represent the linear projection, which map the spatial object prototypes and textual embedding into query and key, respectively.
To facilitate the memory-efficient video understanding, we condense $\mathbf{E}$ to a single token using average pooling. Combining global content token, we extract compact embedding,
$\mathbf{U}_{t,b_{t}=0}=\mathcal{F}_{\text{coarse}}\left(\mathrm{Avg}(\mathbf{E})||\mathrm{Avg}(\mathbf{G}_{t})\right),$ (9)
where $\mathcal{F}_{\text{coarse}}$ is a simple MLP network, and $\mathbf{U}_{t,b_{t}=0}\in\mathbb{R}^{2d^{\prime}}$. Due to that $\mathbf{U}_{t,b_{t}=0}$ only has two tokens, it enables smooth temporal transition and improves the scene consistency for consecutive frames.
Cooperative Encoding.
Given embeddings $\mathbf{U}_{t,b_{t}=1}$ and $\mathbf{U}_{t,b_{t}=0}$ from Cones and Rods, respectively, we combine them to obtain the dynamic embedding of the $t$-th frame,
$\mathbf{O}_{t}=b_{t}\cdot(\mathbf{U}_{t,b_{t}=1}||\mathbf{U}_{t,b_{t}=0})+(1-b_{t})\cdot\mathbf{U}_{t,b_{t}=0}.$ (10)
The video embedding $\mathbf{O}=\{\mathbf{O}_{t}\}_{t=1}^{T}$ and the text embedding $\mathbf{Q}$ are translated into the language space in token format, which is used to generate response from LLMs.
3.4 Training Strategy
In this work, we adopt a two-stage training scheme following previous work [28].
Stage1: Vision-Language Alignment. In the first stage, we pre-train our dynamic cooperative network while freezing both the visual encoder and LLM. It is noteworthy that we only preserve the parameter of projector $\mathcal{F}_{\text{fine}}(\cdot)$ and $\mathcal{F}_{\text{coarse}}(\cdot)$ as the initialization in the second stage. Freezing LLM in the first stage is crucial to effectively align the representation space between video content and language without sacrificing any discernible performance of LLMs.
Stage2: Instruction Tuning. After the first stage, the model possesses the ability of understanding the image within the language space, but fails to flexibly generate the reasonable and coherent linguistic responses. Therefore, in the second stage, we fully fine-tune the LLM and overall parameters in DPE module and CCE module on a instruction-following dataset. This dataset is a composite of pure text QA pairs, single- or multi-turn image QA pairs, and video QA pairs presented in a conversational format. In terms of instruction formulation, different formats are adopted for different kinds of input, and input $\mathrm{\left\langle prompt\right\rangle}$ vary with datasets. Meanwhile, the image token $\mathrm{\left\langle image\right\rangle}$ denotes the placeholder of image or videos, which is randomly inserted at the beginning or end of user prompt or question when training.
4 Experiments
4.1 Experimental Setup
Implementation Details. We use the pre-trained ViT-G/14 from EVA-CLIP [12] as the visual encoder to extract the features of each frame in video, and it can be further changed to other clip-based video encoders. We use pre-trained Qformer weight from InstructBLIP [9] as the textual encoder. Besides, we adopt the Vicuna-7B model [7] as our foundational LLM. Our model is trained using 8 $\times$ NVIDIA A100 80G GPUs.
Training Datasets. We leverage image-video joint training following most of works to enhance the multi-modality understanding of LLMs. Specifically, we leverage the image-to-text dataset LLaVA-filter-CC3M [47] image-caption pairs for the first stage training following LLaVA-VID [28], and LLaVA-665K [13, 20, 23, 24, 32, 38, 39, 46, 48] image QA pairs and ScienceQA [36] for the second stage training, respectively. For video-to-text dataset preparation, we use WebVid-2.5M [2] video-caption pairs for the first stage, and a subset from VideoChat2 for the second stage, including VideoChatGPT-100K [37], WebVid-10M-QA [2], NExT-QA [3], and CLEVRER [62]. And all the samples are formulated as the uniform input format as LLaMA-VID [28].
Table 1: Performance comparisons on zero-shot QA benchmark, including MSVD-QA [58], MSRVTT-QA [59], and ANet-QA [4]. We empirically observe that the default version of GPT-3.5-Turbo would significantly impact evaluation performance. Thus, we also report the possible GPT-3.5 versions for evaluation.
Methods Size MSVD-QA MSRVTT-QA ANet-QA
Acc Score Acc Score Acc Score
VideoLLaMA [63] 7B 51.6 2.5 29.6 1.8 12.4 1.1
LLaMA-Adapter [64] 7B 54.9 3.1 43.8 2.7 34.2 2.7
VideoChat [25] 7B 56.3 2.8 45.0 2.5 26.5 2.2
VideoChatGPT [37] 7B 64.9 3.3 49.3 2.8 35.2 2.7
BT-Adapter [CVPR 24] [34] 7B 67.5 3.7 57.0 3.2 45.7 3.2
Chat-UniVi [CVPR 24] [21] 7B 65.0 3.6 54.6 3.1 45.8 3.2
LLaMA-VID [ECCV 24] [28] 7B 69.7 3.7 57.7 3.2 47.4 3.3
LLaMA-VID [ECCV 24] [28] 13B 70.0 3.7 58.9 3.3 47.5 3.3
VideoChat2 [ECCV 24] [26] 7B 70.0 3.9 54.1 3.3 49.1 3.3
[…]
Table 2: Performance comparisons on VCG-Bench. $\dagger$ represents the version that first fine-tuned on all the dataset, and further post-tuned on VideoChatGPT-100K [37] with a smaller learning rate.
Methods Size CI DO CU TU CO Avg.
VideoLLaMA [63] 7B 1.96 2.18 2.16 1.82 1.79 1.98
LLaMA-Adapter [CVPR 23] [64] 7B 2.03 2.32 2.30 1.98 2.15 2.16
VideoChat [25] 7B 2.23 2.50 2.53 1.94 2.24 2.29
VideoChatGPT [ACL 24] [37] 7B 2.40 2.52 2.62 1.98 2.37 2.38
BT-Adapter [CVPR 24] [34] 7B 2.68 2.69 3.27 2.34 2.46 2.69
[…]
LLaMA-VID [ECCV 24] [28] 7B 2.96 3.00 3.53 2.46 2.51 2.89
VideoChat2 [CVPR 24] [26] 7B 3.02 2.88 3.51 2.66 2.81 2.98
[…]
DynFocus 7B 3.12 3.11 3.68 2.57 2.74 3.05
[…]
Methods Input Holistic Single Detail Multi Detail M-Avg G-Avg
TR AR VS NQA ER PQA SSC AO AC
Short Video MLLMs
VideoChat [25] 16 fr 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
Video-ChatGPT [ACL 24] [37] 100 fr 26.9 24.0 2.31 40.3 42.0 29.9 5.48 25.1 31.1 31.3 3.90
Video-LLaMA2 [6] 16 fr 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
[…]
Long Video MLLMs
[…]
Movie-LLM [51] 1 fps 30.0 29.0 2.88 29.6 24.7 24.1 5.00 20.5 24.8 26.1 3.94
TimeChat [CVPR 24] [44] 96 f 23.1 27.0 2.54 24.5 28.4 25.8 4.29 24.7 32.0 30.9 3.42
LLaMA-VID [ECCV 24] [28] 1 fps 50.8 34.5 3.22 30.1 32.7 32.5 5.22 23.9 27.8 33.2 4.22
MA-LMM [CVPR 24] [16] 1000 f 51.9 35.5 2.12 43.1 38.9 35.8 4.80 25.1 24.3 36.4 3.46
[…]
GPT-4o [41] 0.5 fps 87.4 74.5 4.90 64.8 57.1 65.1 6.69 56.7 46.3 64.6 5.80
4.2 Evaluation on Short Video Understanding
Zero-shot Video-question Answering Performance. In Table 1, we report the results of our DynFocus against a bunch of SOTA methods on three widely-used QA benchmarks: MSVD-QA [5], MSRVTT-QA [60], and ANet-QA [17]. On MSRVTT-QA and MSVD-QA, our model achieves comparable results than published SOTA ST-LLM [35]. For slightly longer video ANet-QA, our method achieves competitive performance using $\sim$25fewer tokens than ST-LLM, exhibiting a balance between accuracy and memory efficiency. Beyond that, we empirically observe that the marginal performance gain on short video dataset gradually decreases as the dataset scale expands during instruction tuning, which can be found in the supplementary material.
VCG-Bench Performance. Table 2 presents the results on VideoChatGPT [37] in terms of Correctness of Information (CI), Detailed Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU) and Consistency (CO). Our DynFocus outperforms existing video MLLMs on CI, DO, and CU. Notably, it substantially surpasses VideoChat2 [26] on CI despite using fewer instructional dataset. This may be attributed to our DPE module, which supports dynamically mitigating the visual nuisance that could hamper factual correctness. ST-LLM shows slight advantages over ours on TU for two possible reasons: (1) it performs the feature alignment between masked input and unmasked video input, which explicitly emphasizes the temporal relationship. (2) The retained tokens of each frame in ST-LLM is more than ours, and more visual details could compensate for temporal clues when handling short videos.
Table 4: Comparisons on VideoMME with short, medium, and long durations, under the settings of “without subtitles” and “with subtitles”. Notably, our method adopts 2242 frame resolution instead of using original resolution. $\dagger$ denotes the model with DPO tuning.
Models Input LLM Size Short (%) Medium (%) Long (%) Overall (%)
w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs
LLaMA-VID [ECCV 24] [28] 1fps 7B - - - - - - 25.9 -
Video-LLaVA [EMNLP 24] [30] 8 fr 7B 45.3 46.1 38.0 40.7 36.2 38.1 39.9 41.6
ST-LLM [ECCV 24] [34] 16 fr 7B 45.7 48.4 36.8 41.4 31.3 36.9 37.9 42.3
[…]
LLaVA-NeXT [65] - 34B 61.7 65.1 50.1 52.2 44.3 47.2 52.0 54.9
VILA-1.5 [31] - 34B 68.1 68.9 58.1 57.4 50.8 52.0 59.0 59.4
Table 5: Performance Comparisons on LV-Bench. Input shows the number of frames each model can process.
Method Size Input ER EU KIR TG Rea Sum Overall
Short Video MLLMs
TimeChat [CVPR 24] [44] 7B 96 fr 21.9 21.7 25.9 22.7 25.0 24.1 22.3
PLLaVA [61] 34B 16 fr 25.0 24.9 26.2 21.4 30.0 25.9 26.1
LLaVA-NeXT [65] 34B 32 fr 30.1 31.2 34.1 31.4 35.0 27.6 32.2
GPT-4o [41] - 10 fr 26.5 23.7 28.3 21.4 28.0 32.8 27.0
Long Video MLLMs
MovieChat [CVPR 24] [49] 7B $\sim$10k fr 21.3 23.1 25.9 22.3 24.0 17.2 22.5
LLaMA-VID [ECCV 24] [28] 13B $\sim$10k fr 25.4 21.7 23.4 26.4 26.5 17.2 23.9
LWM [33] 7B $\sim$4k fr 24.7 24.8 26.5 28.6 30.5 22.4 25.5
Gemini 1.5 Pro [43] - 3.6k fr 32.1 30.9 39.3 31.8 27.0 32.8 33.1
[…]
Figure 4: We showcase the filtered event prototypes focused by DPE module on LV-Bench. To save space, we only showcase the prototype with top-6 score sequentially. The figure at the right-bottom corner illustrates the learned score distribution on the event prototype candidates ($L$=25) obtained by DPC-KNN.
4.3 Evaluation on Long Video Understanding
To demonstrate the advantage of our dynamic cooperative setting, we conduct experiment on three newly released long-term video benchmark. The detailed description for each benchmark are elaborated in supplementary material. MLVU-Bench performance. The evaluation task of MLVU can be categorized into three types: (1) holistic LVU, which requires to make use of global perspectives from the entire video; (2) single-detail LVU, which needs to pinpoint one critical details in a haystack; (3) multi-detail LVU, which calls for the joint utilization of multiple detailed plots within the long video to collaborately infer the answer.
The performance of individual task and the average performance of multi-choice task (M-Avg, within 0-100%) and generation task (G-Avg, within 0.0-10.0) are both reported in Table 3. We have following observations: (1) our DynFocus surpasses all the open-sourced video MLLMs with a clear-cut performance gain on M-Avg and G-Avg, and it nearly consistently ranks top-2 position on individual tasks. (2) For TR, AR, and VS tasks that require an thorough understanding of entire video, our method achieves the best. We attribute this to our dynamic cooperative network’s ability to balance intricate spatial details with broader temporal perception without introducing external visual nuisance. (3) Most approaches find AO and AC tasks challenging due to their sensitivity to the temporal clues, which requires recalling multiple nuanced details from lengthy videos. Although not being further fine-tuned on long-term video dataset like MovieChat, our model still performs competitively. (4) However, our method struggles with ER task that needs ego-based perspectives, likely due to the requirement for ego-centric dataset like EgoQA [10] in VideoChat2.
LV-Bench Performance. We assess six core capabilities of our model on LV-Bench: Temporal Grounding (TG), Summarization (Sum), Reasoning (Rea), Entity Recognition (ER), Event Understanding (EU), and Key Information Retrieval (KIR). The average duration of each video exceeds 1 hour. Following [53], we select several publicly evaluated methods as baselines, with results shown in Table 5. Interestingly, some methods that excel on short videos perform almost randomly in answer selection. Remarkably, our DynFocus achieves the best of 31.5% among all the open-sourced 7B models, even outperforming PLLaVA [26] with 34B parameters.
((a))
((b))
Figure 5: (a) and (b) illustrate the performance with different number of event prototypes and different ratio of filtered event prototypes, respectively.
VideoMME Benchmark Performance. VideoMME benchmark spans across three kinds of durations, VideoMME-S ($\sim$1.3min), VideoMME-M ($\sim$8.5min), and VideoMME-L ($\sim$0.7h). Most videos include both subtitles and audios, which helps us investigate the performance gain from additional information sources. Figure 4 compares our results with other representative video MLLMs. Notably, our DynFocus consistently achieves impressive advantage across different lengths of video with subtitles and without subtitles. Specifically, it exhibits the overwhelming advantage over SOTA ST-LLM and VideoChat2. Remarkably, the version of DynFocus without subtitles reaches an overall accuracy of 44.1%, still surpassing ST-LLM with subtitles by 1.8%.
4.4 Evaluation on Video Hallucination
We report the results about addressing video hallucination on VideoHallucer [54] and give further analysis in the supplementary materials due to the space limitation.
4.5 Component-wise Analysis
[…]
$\left|\mathbf{U}_{b_{t}=0}\right|$ $\left|\mathbf{U}_{b_{t}=1}\right|$ MSVD-QA ANet-QA VCG-Bench
Acc Score Acc Score Score
0 40 63.7 3.5 41.4 3.2 2.57
0 256 65.6 3.5 42.1 3.2 2.65
2 256 68.4 3.7 44.3 3.4 2.85
2 2 62.0 3.5 40.5 3.2 2.38
2 0 58.2 3.3 38.6 2.9 2.21
2 40 67.9 3.7 43.1 3.3 2.81
Figure 6: Token number comparison with different methods on different benchmark datasets. We calculate their token number using their released code snippet regarding loading video.
The Effect of Number of Initial Event Prototypes. As depicted in Figure 5a, we observe that increasing $L$ (10 $\sim$ 25) brings a consistent gain in the overall accuracy. It indicates that sufficient number of prototypes could divide the video into more fine-grained events, which would offer more abundant visual clues for accurate question answering. However, when $L$ increases greater than 25, the performance begins to drop. This phenomenon can be explained by that increasing prototypes would hamper the intrinsic temporal structure as well as consistency.
The Effect of Dynamic Selection. As shown in Figure 5b, a similar pattern can be observed by varying the ratio $K/L$, where $L$ and $K$ represent the number of event prototype candidates obtained by DPC-KNN clustering and filtered ones, respectively. The smaller ratio of filtered prototypes may be not enough to cover all useful visual clues, whereas the larger ratio still maintain much non-essential visual nuisance, thereby disturbing content understanding.
Visualization of Focused Frame by DPE module. As explained in DPE module, the higher score in Figure 4 indicates greater contribution to question answering. Taking the first case as an example, the question asks about the occurred events and their corresponding sequence. Although these frames with great contribution encoded by Cones could provide detailed visual semantics like sea, waterfall, and aurora, they may not offer sufficient temporal clues to determine the order of events. This information can be encapsulated by its complementary part, akin to Rods, which provides broader receptive field for capturing motion.
The Effect of Cooperation between Cones and Rods. We introduce several variants to validate the benefit of cooperation between Cones and Rods. As reported in Table 6, $\left|\mathbf{U}_{b_{t}=0}\right|=0$ indicates that we discard the tokens encoded by Rods. We observe significant performance drop 1.7% on MSVD-QA compared with our full model. The similar pattern can be observed when completely dropping the tokens encoded by Cones. Moreover, although the model exhibits the best results without token compression for those important frames, it still encounter the scalability issues when extending to long-term videos, struggling to balance memory efficiency and accuracy. Figure 6 shows the comparison of total token usage, demonstrating superior advantages over existing methods.
The Effect of Compact Encoding in CCE. As shown in Table 7, we introduce several variants to assess the impact of fusion strategies between event prototypes $\mathbf{h}_{t}$ and spatial multi-grained prototypes $\mathbf{G}_{t}$ on model performance. Although direct concatenation uses slightly more tokens compared to cross-attention, it offers performance advantages with greater parameter efficiency, making it our paramount choice.
Model Variants MSVD-QA VCG-Bench
Acc Score Score
Cross-attention (Soft) 64.74 3.61 2.56
Concat. 66.20$\uparrow$ 3.67$\uparrow$ 2.66$\uparrow$
Concat. + Multi-grained 67.90$\uparrow$ 3.72$\uparrow$ 2.81$\uparrow$
Table 7: Effects of different components in CCE module. Concat. is the concatenation operation.
5 Conclusion
In this paper, we develop a dynamic cooperative network for memory-efficient encoding. We experimentally delve into the network behavior and find that dynamic encoding could simultaneously achieves fine spatial visual appearance understanding and coarse temporal dynamics perception using affordable tokens, striking a balance between answering accuracy and memory efficiency. Our model achieves superior performance with substantially few tokens on both short and long video benchmarks. Moreover, our model also demonstrates the great potential on addressing video hallucination.